OpenStudy (anonymous):

Solve for x: 7 over 8 minus 1 over x equals 3 over 4

OpenStudy (anonymous):

Is it: \[\frac78 - \frac1x= \frac34\]?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

\[\frac{7}{8} - \frac{1}{x} = \frac{3}{4}\] subtract 7/8 from both sides then \[-1/x = -1/8\] x = 8